# Expanded Generalized Needlet Internal Linear Combination (eGNILC) Framework for the 21-cm Foreground Removal

### v1

@@ 7 Acknowledgements
We thank the useful comment from Prof. C. Dickinson. YZM is supported by the National Research Foundation of South Africa under Grants No. 150580, No. 159044, No. ERC23040389081 and No. CHN22111069370. This work was part of the research programme “New Insights into Astrophysics and Cosmology with Theoretical Models Confronting Observational Data” of the National Institute for Theoretical and Computational Sciences of South Africa.
@@ Appendix A Needlet analysis
where $P_{\ell}$ functions are the Legendre polynomials. With Eqs. (A1) and (A2), isotropic wavelet analysis can be implemented either in the spatial domain or in the harmonic domain. Here, we choose the latter by multiplying the harmonic coefficients of the field of interest $X$ with a spectral window $h_{\ell}$. We call exact reconstruction condition that ensures $\sum_{j\in\mathcal{J}}\Psi^{(j)}=\textbf{Id}$, or equivalently
@@ Appendix B Calculation of the eGNILC bias and AIC with $m$-modes of foreground and $N_{\rm p}$ pixel data
where in the third equality, we have used the property of the expectation value ${\rm E}\left(\bm{s}^{T}_{p}\bm{s}_{q}\right)=\bm{R}_{\bm{s}}\delta_{pq}$. We have also used the matrix property (see also Hanson & Lewis 2009)
@@
where in the fourth equality, we have used the property that $\bm{S}$ and $\bm{R}_{\bm{x}}$ matrices are of the dimensions $n_{\rm ch}\times(n_{\rm ch}-m)$ and $n_{\rm ch}\times n_{\rm ch}$ respectively (Eqs. (37) and (40)).

### v2

@@ 7 Acknowledgements
We thank the useful comment from Prof. C. Dickinson. Y.Z.M. is supported by the National Research Foundation of South Africa under grant Nos. 150580, 159044, ERC23040389081 and CHN22111069370. This work was part of the research programme “New Insights into Astrophysics and Cosmology with Theoretical Models Confronting Observational Data” of the National Institute for Theoretical and Computational Sciences of South Africa.
@@ Appendix A Needlet analysis
where $P_{\ell}$ functions are the Legendre polynomials. With Eqs. (A1) and (A2), isotropic wavelet analysis can be implemented either in the spatial domain or in the harmonic domain. Here, we choose the latter by multiplying the harmonic coefficients of the field of interest $X$ with a spectral window $h_{\ell}$. We call this the exact reconstruction condition that ensures $\sum_{j\in\mathcal{J}}\Psi^{(j)}=\textbf{Id}$, or equivalently
@@ Appendix B Calculation of the eGNILC Bias and AIC with $m$-modes of Foreground and $N_{\rm p}$ Pixel Data
where, in the third equality, we have used the property of the expectation value ${\rm E}\left(\bm{s}^{T}_{p}\bm{s}_{q}\right)=\bm{R}_{\bm{s}}\delta_{pq}$. We have also used the matrix property (see also Hanson & Lewis 2009)
@@
where, in the fourth equality, we have used the property that $\bm{S}$ and $\bm{R}_{\bm{x}}$ matrices are of the dimensions $n_{\rm ch}\times(n_{\rm ch}-m)$ and $n_{\rm ch}\times n_{\rm ch}$ respectively (Eqs. (37) and (40)).